OpenStudy (anonymous):

3x-2y=18 Need x and y and slope

OpenStudy (anonymous):

I'm assuming you want the x-intercept, y-intercept, and slope. The easiest way to do this is to solve the equation for y (which will put the equation into slope-intercept form, y = mx + b). 3x - 2y = 18 2y = 3x - 18 y = (3/2)x -9 The slope is then 3/2, and the y-intercept is at (0,-9). Finally, we plug in 0 for y in the original equation and solve for x to get the x-intercept. Doing so gives us: 3x -2(0) = 18 3x = 18 x = 6 So the x-intercept is (6,0).